Circular No. 7018

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998eb IN NGC 1961 S. Jha, P. Garnavich, P. Challis, and R. Kirshner, Harvard-Smithson-

ian Center for Astrophysics, report that a spectrum of SN 1998eb, taken by P. Berlind on Sept. 19.5 UT with the F. L. Whipple Observatory 1.5m telescope (+ FAST spectrograph), reveals it to be a type-Ia supernova, three to four weeks after maximum light. The recession velocity of the host galaxy is 3900 km/s, measured via a narrow emission line of Hα.

A. V. Filippenko and D. C. Leonard, University of California at Berkeley, report that a CCD spectrum (range 340–1000 nm) obtained on Sept. 20 with the Shane 3-m reflector at Lick Observatory shows that SN 1998eb (cf. IAUC 7016) is a type-Ia supernova, a few weeks past maximum brightness.

SUPERNOVA 1998dw IN ANONYMOUS GALAXY Filippenko and Leonard also report that a CCD spectrum obtained on

Sept. 21 with the Shane 3-m reflector shows that SN 1998dw (cf. IAUC 7007) is a type-Ia supernova, a few weeks past maximum brightness.

NOVAE IN NGC 221 AND M31 Filippenko and Leonard confirm two of the recent novae in NGC 221

(cf. IAUC 7004) and M31 (cf. IAUC 7015). Prominent Balmer emission lines are visible in the spectra (range 430–690 nm), and the spectrum of the latter is of sufficient quality to reveal Fe II emission as well.

COMET C/1995 O1 (HALE-BOPP)

Ephemeris extension to IAUC 6942 (elements MPC 32410):

1998TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Sept. 24 7^{h}26.^{m}91 −61^{o}58.^{′}9 6.580 6.482 80.^{o}0 8.^{o}8 10.2
Oct. 4 7 29.34 −63 45.3 6.657 6.570 80.7 8.6 10.3
14 7 29.69 −65 31.7 6.735 6.657 81.3 8.5 10.4
24 7 27.54 −67 15.5 6.813 6.743 81.8 8.4 10.5
Nov. 3 7 22.44 −68 54.0 6.894 6.829 82.1 8.3 10.5
13 7 14.0 −70 23.8 6.975 6.915 82.4 8.2 10.6
23 7 02.0 −71 41.4 7.057 7.000 82.6 8.0 10.7

Visual m_{1} estimates: July 29.77 UT, 9.1 (D. A. J. Seargent, The Entrance, N.S.W., 10×50 binoculars); Sept. 1.72, 9.4 (Seargent); 16.62, 10.2 (M. Mattiazzo, Wallaroo, South Australia, 0.20-m reflector).

1998 September 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green